# Machine learning classification of black holes in the mass–spin diagram

### v2

@@ Acknowledgements.
The Authors thank Nicole Ndegwa for support, inspiration, and perspicacious suggestions for visualizations. We also thank Davide Gerosa for helpful feedback on machine learning methods, Michael Kesden for inspiring early comments, and Austin MacMaster for insightful discussions. N.S. and S.S.H. are supported by the Natural Sciences and Engineering Research Council of Canada through the Canada Research Chairs and Discovery Grants programs. This research was enabled by use of the GREX cluster (https://um-grex.github.io/grex-docs/friends/alliancecan/) and the Digital Research Alliance of Canada (alliancecan.ca). Simulations in this paper made use of PyTorch [40], Scikit-learn [41], SciPy [113], NumPy [136], Pandas [137, 138], h5py [139], matplotlib [140], and the COMPAS rapid binary population synthesis code (version 02.42.01) which is freely available at http://github.com/TeamCOMPAS/COMPAS.

### v4

@@ Acknowledgements.
The Authors thank Nicole Ndegwa for support, inspiration, and perspicacious suggestions for visualizations. We also thank Davide Gerosa for helpful feedback on machine learning methods, Michael Kesden for inspiring early comments, Austin MacMaster for insightful discussions, and the anonymous referee for constructive recommendations. N.S. and S.S.H. are supported by the Natural Sciences and Engineering Research Council of Canada through the Canada Research Chairs and Discovery Grants programs. This research was enabled by use of the GREX cluster (https://um-grex.github.io/grex-docs/friends/alliancecan/) and the Digital Research Alliance of Canada (alliancecan.ca). Simulations in this paper made use of PyTorch [44], Scikit-learn [45], SciPy [116], NumPy [140], Pandas [141, 142], h5py [143], matplotlib [144], and the COMPAS rapid binary population synthesis code (version 02.42.01) which is freely available at http://github.com/TeamCOMPAS/COMPAS.